OpenStudy (anonymous):

A simple lever is used to lift a heavy load. When a 60-N force pushes one end of the lever down 1.2 m, the load rises 0.2 m. Show that the weight of the load is 360 N How do I solve

OpenStudy (anonymous):

You are doing some work to lift the load i.e you are increasing the potential energy of the load by doing work. So Work done = PE of the load Does that help ?

OpenStudy (anonymous):

i really do not understand how to do the formulas or how to work them out. this is the first week of class my instructor is no good and a required coarse

OpenStudy (anonymous):

60*1.2/0.2?

OpenStudy (anonymous):

or 360*1.2/60-0.2?

OpenStudy (anonymous):

Work done = force * displacement (in the direction of force) = 60N * 1.2 m PE of the load = weight * height = W* 0.2m So W*0.2 = 60*1.2 (Omitting units for the time being) W = 60*1.2/0.2

OpenStudy (anonymous):

i see thank you